A box contains 72 coins, only dimes and nickels. The amount of money in the box is $5.50. How many dimes and how many buckles are in the box?
Answer by ewatrrr(24785)


Hi
Let x represent the number of Dimes
10(x) + 5(72-x) = 550 |CENTS makes sense
5x = 550 - 360 = 190
x = 38 Dines and 34 Nickels
380 + 170 = 550 Checks!!!